# ScholaWrite: A Dataset of End-to-End Scholarly Writing Process

### v3

@@ Abstract
Writing is a cognitively demanding task involving continuous decision-making, heavy use of working memory, and frequent switching between multiple activities. Scholarly writing is particularly complex as it requires authors to coordinate many pieces of multiform knowledge. To fully understand writers’ cognitive thought process, one should fully decode the end-to-end writing data (from individual ideas to final manuscript) and understand their complex cognitive mechanisms in scholarly writing. We introduce ScholaWrite dataset, a first-of-its-kind keystroke corpus of an end-to-end scholarly writing process for complete manuscripts, with thorough annotations of cognitive writing intentions behind each keystroke. Our dataset includes LaTeX-based keystroke data from five preprints with nearly 62K total text changes and annotations across 4 months of paper writing. ScholaWrite shows promising usability and applications (e.g., iterative self-writing), demonstrating the importance of collection of end-to-end writing data, rather than the final manuscript, for the development of future writing assistants to support the cognitive thinking process of scientists. Our de-identified data examples and code are available on our project page111https://minnesotanlp.github.io/scholawrite/.
@@
Linghe Wang††thanks: Equal contribution. Minhwa Lee Ross Volkov Luan Tuyen Chau Dongyeop Kang University of Minnesota {wang9257,lee03533,volko032,chau0139,dongyeop}@umn.edu
@@ 1 Introduction
Scholarly writing requires researchers to produce texts that concisely and precisely present novel findings and to follow a systematized structure and style by their targeted venue Jourdan et al. (2023); Kallestinova (2011); Bourekkache (2022). To address this need, researchers have leveraged large language models (LLMs) to develop intelligent support systems in several writing tasks, such as revision process Du et al. (2022b); Kim et al. (2022) or feedback generation Liang et al. (2024); Lu et al. (2024a). However, a critical question arises: Are LLMs able to truly understand and generate writing as human scientists do?
LLMs are generally trained to progress autoregressively, generating text from left to right. In contrast, human writing typically involves multiple iterations of complex and non-linear cognitive actions to refine the main message Flower and Hayes (1981); Koo et al. (2023), as illustrated in Figure 1. This fundamental cognitive gap between human writing and language models necessitates a deeper understanding of human writing processes and the development of cognitively-aligned writing assistance tools, moving beyond basic auto-completion.
Although much literature has examined the cognitive processes used in general writing tasks Lindgren and Sullivan (2003); Hayes and Berninger (2014); Zhu et al. (2023); Wengelin and Johansson (2023), the specific cognitive patterns in scholarly writing remain understudied. Our work observes the end-to-end cognitive process of scholarly writing, inspired by the method of keystroke collections that have been a major methodology to observe individual writing processes Baaijen et al. (2012); Latif (2008); Leijten and Van Waes (2013) in cognitive science or a few recent works in NLP at a small scale Koo et al. (2023); Velentzas et al. (2024).
We believe our work (ScholaWrite) is the first attempt to present a keystroke corpus of scholarly writing with annotations of cognitive writing intentions, which were collected over multiple months and produced by early-career researchers. Furthermore, we developed novel data collection and annotation systems, which improves data accessibility to the public and enables real-time collection of LaTeX-based keystrokes from multiple researchers. We also present a comprehensive taxonomy of cognitive writing processes specific to the scholarly writing domain. Throughout several experiments with LLMs, ScholaWrite shows promising usability as a resource for enabling LLMs to understand the human cognitive process of scholarly writing and provide writing suggestions that are aware of writers’ cognitive behaviors.
Contributions
We present ScholaWrite, a curated dataset of nearly 62K LaTeX-based keystrokes that were turned into publications in the computer science domain, annotated by experts in linguistics and computer science. We develop a taxonomy of cognitive process of scholarly writing, providing an overall understanding of how scholars tend to develop writing manuscripts. Experiment results showed that the Llama-8B model fine-tuned on the ScholaWrite dataset achieved the high linguistic quality of the final writing.
@@ 2 Related Work
2.1 Cognitive Theory of Writing Process
@@
Research on human writing shifted from analyzing final written products to examining cognitive processes across various writing phases Diederich (1974); Krapels (1990); MacArthur and Graham (2016). Building on this process-oriented approach, Flower and Hayes (1981)’s cognitive writing theory outlines three key sub-processes: (1) planning, (2) translating, and (3) reviewing. These interconnected and non-linear stages of human writing inform our work, where we expand this model to a more granular taxonomy of cognitive writing patterns in scholarly communication.
Koo et al. (2023) introduced a taxonomy of scholarly writing process based on keystroke data from short research plans written in 30-minute sessions. Expanding on this, we perform a larger-scale keystroke collection over months, whose final products culminated in research publications. Expert-reviewed and grounded in prior literature, our taxonomy encompasses end-to-end cognitive trajectories of scholarly writing.
2.2 Keystroke Loggers for Scholarly Writing
Keystroke logging tools (e.g., Inputlog) allow researchers to observe digital writing without disrupting the writing process Chan (2017); Johansson et al. (2010); Leijten and Van Waes (2013); Lindgren and Sullivan (2019). However, they are often restricted to a closed ecosystem, such as MS Word, and thus less accessible for scientific communities who use LaTeX. Moreover, they are not well-suited for collecting data from extended writing sessions.
Current studies rarely examined scholarly communication that uses LaTeX, particularly for scientific writing. To address this gap, we developed systems that securely collect real-time keystrokes over months from Overleaf, a widely-used online LaTeX editor, while ensuring privacy. The system workflow enables span-level annotation of writing intentions, offering a natural, uninterrupted environment for studying cognitive writing activities over long periods and across multiple sections.
2.3 Datasets for Scholarly Writing
@@
Previous work primarily focused on constructing datasets to analyze scholarly writing processes, which vary in content and scale. Publicly available datasets tend to track linguistic style changes or grammatical edits during revision Du et al. (2022b); Jiang et al. (2022); Ito et al. (2019); Mita et al. (2022), while others capture edits based on feedback and peer review D’Arcy et al. (2024); Jourdan et al. (2024); Kuznetsov et al. (2022) or citation generation Kobayashi et al. (2022); Narimatsu et al. (2021). Also, most focus on specific sections of papers (e.g., abstracts, introductions) Du et al. (2022b); Mita et al. (2022). In contrast, our dataset covers all cognitive phases of writing, from inception to final manuscript, over an extended period.
Furthermore, those existing corpora often compare multiple versions of final manuscripts from preprint databases Jiang et al. (2022); Du et al. (2022b); D’Arcy et al. (2024); Jourdan et al. (2024), missing how those manuscripts evolved Jourdan et al. (2023). To address this, we build a keystroke-based corpus that captures the progression of publications in real-time (See Figure 2).
3 Dataset Construction
To collect authentic writing data from writers’ natural thoughts, we developed a Chrome extension that records keystroke data in the background. After the study, an in-depth annotation process is conducted to interpret the intent behind each keystroke. We describe the detailed process of our data collection, annotation system design, and taxonomy creation in the following sections.
@@ 3.1 Participant Recruitment
We recruited 10 graduate students222All participants attend a R1 university in US and are proficient in English. Appendix A.1 details the recruitment process. from the computer science department who were actively preparing academic manuscripts in Overleaf LaTeX editor for submission to peer-reviewed conferences in AI and NLP. The data is collected from November 2023 to February 2024, totaling up to 4 months. Our study has been approved by the IRB institution of the authors. Please refer to Appendix A.1 for more detailed descriptions.
3.2 Chrome Extension for Keystroke Collection in Overleaf
We designed and implemented a Chrome extension (Appendix Figure 7), which enables the real-time collection of keystroke trajectories in the Overleaf platform. Participants can create their account credentials, and after logging into the system, the extension monitors the user’s keystrokes in the background silently without disrupting writing process. See Appendix A.2 for more description of the extension workflow.
@@ 3.2 Chrome Extension for Keystroke Collection in Overleaf
Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing keywords or notes (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
@@
[…]
Coherence | Logically link (1) any of the two or multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent as a whole. | “Each comment was annotated by three different annotators , which and we achieved high inter-annotator agreement.” | 3.3% | |
@@
Structural | Improve the flow of information by modifying the location of texts and objects. | “We calculate Pearson’s correlation for human alignment to compare the alignment between lexicon-based preferences and humans’ original preferences. First, we calculated the score from each human participant to compare the alignment between lexicon-based preferences and humans’ original preferences.” | 3.7% | |
@@
3.3 Writing Intention Annotation
To decode the nuanced intentions behind each writing action, we annotate each keystroke collected from the Chrome extension, using a specialized interface. This process is crucial for developing a detailed taxonomy of the end-to-end scholarly writing process, which serves as the basis for annotating the collected keystrokes. To perform annotations, we processed those raw keystroke entries by file name, type of writing actions, an array of differences between two subsequent texts, and line number in an Overleaf editor. Please see Appendix B for further annotation details.
Annotation Interface
We developed a novel interface for keystroke annotation, enabling visualization by time, LaTeX file, and author (Appendix Figure 8). Annotators can navigate the keystroke timeline, identify a span of related edits (e.g., outlining or grammar correction), and assign a pre-defined intention label from the dropdown menu.
@@ 3.3 Writing Intention Annotation
Our taxonomy (Table 1) builds on prior literature Flower and Hayes (1981); Du et al. (2022b); Koo et al. (2023). During annotations, we have identified 15 types of scholarly writing intentions from the five projects and further grouped them into three high-level categories according to Flower and Hayes (1981) - Planning, Implementation, and Revision.
Following Pustejovsky et al. (2017), we performed an iterative open coding for the collected keystrokes from the two Overleaf projects. Using the annotation interface, each annotator reviewed the first few writing keystrokes and identified the high-level processes. Next, the annotators (1) determine the span of keystrokes that form a meaningful unit together (e.g., a clause or sentence) within the high-level process and (2) decide the underlying intention of edits. The annotators coded these spans independently, resolving conflicts through multiple in-depth discussions. Finally, a cognitive linguist participated in all discussions, refining the definition and scope of each writing intention.
To ensure taxonomy validity, we followed two standard practices Nickerson et al. (2013); Kundisch et al. (2021): (1) mutual exclusivity and (2) collective comprehensiveness. Ambiguous keystroke spans were flagged and refined through iterative discussions to sharpen category boundaries. After finalizing the codebook and taxonomy, we updated intention annotations333The updated list appears in annotation interface. for the two selected projects and applied the taxonomy to the rest. The final inter-annotator agreement was (weighted F1 score in a multi-label setting for a 1K-keystroke sample444Unlike Kappa, which assumes categorical agreement and struggles with multi-label classification, weighted F-1 scores better handle class imbalances and partial agreements.). See Appendix B for full annotation details.
4 ScholaWrite: A Dataset of Cognitive Process of Scholarly Writing
Summary
Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with annotated intentions, for five Overleaf projects whose final product turned into arXiv preprints. The dataset consists of 61,504 arrays of keystroke changes with fine-grained annotations of writing intentions. Appendix Table 6 shows the overall statistics of the dataset.
Writing Intention Distributions
The writing intention distributions is illustrated in Figure 3(c). During the planning stage, idea generation is the most frequent intention. Text production dominates the implementation stage, and clarity becomes the primary focus during the revision stage. Since we defined the boundary between implementation and revision as the point where meaning changes occur, the high frequency of text production suggests that authors often continue revising their texts by introducing new topics into the documents. Appendix Table 8 shows the label distribution per project.
Label | Subsequent label | Probability | |
Idea Generation | Text Production | 0.52 | |
Idea Organization | Idea Generation | 0.34 | |
Citation Integration | Text Production | 0.37 | |
Cross-reference | Text Production | 0.36 | |
Clarity | Text Production | 0.35 | |
Coherence | Text Production | 0.34 | |
Scientific Accuracy | Text Production | 0.34 |
Flows Between Writing Intentions
We analyze the flow of writing intentions most likely to follow each other throughout the writing process, as presented in Table 2. Many intentions proceed to text production, where the idea generation has the highest probability of feeding into text production. Figure 3(a) also shows a similar pattern of flows between text productions and other intentions. Also, in Table 2 the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims. Please see Appendix Figure 13 for the intention flows of each of the five projects.
Time-series Distributions of Writing Intentions
We calculated the average Wasserstein distance between each intention distribution and uniform distribution to assess how evenly intentions occur throughout the writing process. As shown in Figure 3(d), text production had the smallest distance, while scientific accuracy had the largest. This suggests that text production is spread throughout the writing, whereas scientific accuracy is concentrated in specific phases. A similar pattern appears in Figure 3(e) , where text production is evenly distributed over time, and scientific accuracy peaks in the middle-to-late of the writing process. Figure 3(b) demonstrates a similar pattern that text production of the implementation stage spreads evenly over time. Types of clarity and structural revisions tend to occur after the medium steps of writing. Please see Appendix Figures 9 to 12 for each of the five projects.
5 Applications for Writing Assistance
We envision ScholaWrite as a valuable resource for training language models and improving future writing assistants for scholarly writing. To evaluate its usability, we conducted experiments training LLMs to mimic the complex, non-linear writing processes of human scholars. Specifically, we aimed to showcase the capabilities of LLMs trained on ScholaWrite in two scenarios:
(1) Predicting an author’s next writing intention (§5.1): The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention. This enables them to offer cognitively-appropriate writing suggestions that align with the writer’s needs.
(2) Iteratively generating scholarly writing actions from scratch (§5.2) (called Iterative Self-Writing), mirroring the human writing process: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
5.1 Predicting Next Writing Intention
Setup & Metrics
This task (represented by the inner box “Prediction” in Figure 4) takes the “before-text” from a keystroke pair (i.e., Listing LABEL:table:single-entry) and context prompt, and predicts the writing intention to apply for the subsequent actions.
We use BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and Llama3.1-8B-Instruct Dubey et al. (2024) as baselines, fine-tuning each on the ScholaWrite training set. For comparison, we also run GPT-4o OpenAI (2024) on the test set. To evaluate model performance, we used a weighted F-1 score555Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 3(c) and Table 1).. See Appendix E.2 for details.

### v4

@@ Abstract
Writing is a cognitively demanding activity that requires constant decision-making, heavy reliance on working memory, and frequent shifts between tasks of different goals. To build writing assistants that truly align with writers’ cognition, we must capture and decode the complete thought process behind how writers transform ideas into final texts. We present ScholaWrite, the first dataset of end-to-end scholarly writing, tracing the multi-month journey from initial drafts to final manuscripts. We contribute three key advances: (1) a Chrome extension that unobtrusively records keystrokes on Overleaf, enabling the collection of realistic, in-situ writing data; (2) a novel corpus of full scholarly manuscripts, enriched with fine-grained annotations of cognitive writing intentions. The dataset includes LaTeX-based edits from five computer science preprints, capturing nearly 62K text changes over four months; and (3) analyses and insights into the micro-dynamics of scholarly writing, highlighting gaps between human writing processes and the current capabilities of large language models (LLMs) in providing meaningful assistance. ScholaWrite underscores the value of capturing end-to-end writing data to develop future writing assistants that support, not replace, the cognitive work of scientists.
@@
Khanh Chi Le††thanks: Equal Contribution, Linghe Wang∗, Minhwa Lee∗, Ross Volkov, Luan Tuyen Chau, Dongyeop Kang University of Minnesota {le000422, wang9257, lee03533, volko032, chau0139, dongyeop}@umn.edu
@@ 1 Introduction
Scientific writing is one of the most cognitively demanding forms of human communication. Researchers must articulate novel ideas with precision and structure, iteratively refining arguments, evidence, and phrasing over time (Jourdan et al., 2023; Kallestinova, 2011; Bourekkache, 2022). Unlike linear text generation, scholarly writing is a complex cognitive process, writers continually plan, draft, and revise across multiple stages and intentions (see Fig. 1). Yet, despite the growing integration of large language models (LLMs) into research workflows, our understanding of how human scientists actually write remains limited. Without such understanding, it is difficult to build writing assistants that genuinely align with the cognitive processes of their users.
Recent efforts in writing assistance have leveraged LLMs for tasks such as revision or feedback generation (Du et al., 2022b; Liang et al., 2024). However, these models largely operate autoregressively, producing text from left to right, while human writing unfolds through recursive, non-linear cycles of ideation, organization, and refinement (Flower and Hayes, 1981). This fundamental cognitive gap raises a key question: how can we empirically capture and model the end-to-end scholarly writing process so that future LLMs can support, rather than substitute, human cognition?
Existing studies have examined writing processes through revision logs or small-scale keystroke analyses (Leijten and Van Waes, 2013; Koo et al., 2023), but most focus on short writing tasks or version-level comparisons of finalized drafts (Du et al., 2022b; Jiang et al., 2022; D’Arcy et al., 2024). Such approaches miss the fine-grained temporal and cognitive dynamics that drive real-world scholarly writing over extended periods.
To bridge this gap, we present ScholaWrite, the first dataset capturing the end-to-end scholarly writing process, from initial drafts to final manuscripts, through in-situ keystroke logging and cognitive annotation. Our work contributes three main advances:
(1) A reusable tool and taxonomy for cognitive writing capture. We develop a Chrome extension that unobtrusively records keystrokes within Overleaf, the dominant platform for collaborative scientific writing. Paired with a taxonomy of fifteen cognitive writing intentions, this tool enables large-scale, domain-specific capture of authentic writing behaviors (§3).
(2) The ScholaWrite dataset. Using this system, we collect over 61K keystroke-based text changes from five Computer Science preprints written by graduate researchers over four months. Each keystroke is annotated with its underlying cognitive intention (e.g., idea generation, clarity improvement), producing a richly structured corpus for modeling writing cognition (§4).
(3) Empirical insights into human-LLM gaps. Our analyses reveal that scholarly writing is highly non-linear and multi-intentional: more than half of all writing sessions involve three or more intertwined intentions. Writers dedicate most time to producing and revising text for clarity and coherence, while cognitively heavy tasks, such as idea organization or visual structuring, occur in fewer but longer, more focused sessions. Evaluations of current LLMs (GPT-5, Qwen) show they can imitate surface-level edits but fail to predict next intentions or sustain cognitively complex revisions, underscoring the limits of current models as cognitively aligned writing partners (§5 and §6).
@@ 2 Related Work
Cognitive Theory of Writing Process
@@
Research on writing has shifted from analyzing final texts to examining cognitive processes across writing phases (Diederich, 1974; Krapels, 1990; MacArthur and Graham, 2016). Building on this process-oriented approach, Flower and Hayes (1981) outline three core sub-processes: planning, translating, and reviewing. These dynamic, non-linear stages inform our work, where we extend this model into a finer-grained taxonomy of cognitive writing patterns in scholarly communication. Koo et al. (2023) proposed a taxonomy of scholarly writing based on keystroke data from short, 30-minute research-plan sessions. Expanding on this, we collect larger-scale, longitudinal keystroke data spanning months and culminating in published research. Expert-reviewed and grounded in prior theory, our taxonomy captures end-to-end cognitive trajectories of scholarly writing.
Keystroke Loggers for Scholarly Writing
Keystroke logging tools (e.g., Inputlog) enable unobtrusive observation of digital writing (Chan, 2017; Johansson et al., 2010; Leijten and Van Waes, 2013; Lindgren and Sullivan, 2019). Yet, most operate in closed ecosystems like MS Word and are ill-suited for LaTeX-based academic writing or extended sessions. To address this, we built systems that securely record real-time keystrokes from Overleaf—an online LaTeX editor—over months while preserving privacy. This workflow supports span-level annotation of writing intentions, providing a natural setting to study long-term cognitive writing activities across document sections.
Datasets for Scholarly Writing
@@
Publicly available datasets from previous work tend to track linguistic style changes or grammatical edits during revision (Du et al., 2022b; Jiang et al., 2022; Ito et al., 2019; Mita et al., 2022), while others capture edits based on feedback and peer review (D’Arcy et al., 2024; Jourdan et al., 2024; Kuznetsov et al., 2022) or citation generation (Kobayashi et al., 2022; Narimatsu et al., 2021). Also, most focus on specific sections of papers (e.g., abstracts, introductions) (Du et al., 2022b; Mita et al., 2022). In contrast, our dataset covers all cognitive phases of writing over an extended period and all sections. Furthermore, those existing corpora often compare multiple versions of final manuscripts from preprint databases (Jiang et al., 2022; Du et al., 2022b; D’Arcy et al., 2024; Jourdan et al., 2024), missing how those manuscripts evolved (Jourdan et al., 2023). To address this, we build a keystroke-based corpus that captures real-time progression of publications.
3 Data Collection and Intent Annotation
To capture scholarly writing as it naturally unfolds, we developed a system that records the end-to-end writing process, from individual keystrokes to their cognitive interpretation, without interrupting authors’ workflow. This section introduces the design of our Chrome extension for Overleaf-based data collection, the participant recruitment protocol, and the annotation pipeline that underpins our taxonomy of cognitive writing intentions.
3.1 Keystroke Collection in Overleaf
We built a custom Chrome extension (Appendix Figure 8) that unobtrusively logs real-time keystroke trajectories from Overleaf. The extension operates in the background once participants consent and authenticate via unique credentials. Every time a key-up event occurs, the system captures the visible text within the user’s Overleaf editor, compares it with the previous version using Google’s diff_match_patch algorithm, and stores the resulting text differences alongside metadata such as timestamp, file name, and author ID. To protect privacy, only project IDs pre-approved through participant consent were collected; any non-consented Overleaf projects were automatically filtered out. All data were stored on a secure institutional server and anonymized before analysis. Appendix A.2 provides further implementation details.
@@ 3.2 Participant Recruitment
We recruited 10 graduate students in computer science at an R1 university in the United States, each actively writing manuscripts for peer-reviewe AI and NLP conferences using Overleaf. The collection period spanned four months (November 2023–February 2024), capturing writing sessions over the full lifecycle of manuscript preparation. All participants were proficient in English and provided informed consent under IRB-approved protocols (Appendix A.1). Each participant installed the Chrome extension on their personal device and continued their usual writing activities, allowing us to gather naturalistic data without external disruption.
3.3 Writing Intention Annotation
To uncover the cognitive mechanisms underlying scholarly writing, we annotate every collected keystroke with its corresponding writing intention. This annotation process transforms raw writing traces into interpretable cognitive data, forming the foundation of the ScholaWrite taxonomy and subsequent analyses. Each keystroke record contains metadata, such as file name, type of action, text differences between two states, and line number in the Overleaf editor, which allows precise reconstruction of how writers iteratively develop and refine their manuscripts. Please see Appendix B for further annotation details.
Annotation Interface
We developed an interactive annotation interface that visualizes writing activity over time and across authors and files (Figure 9 in Appendix). Annotators can navigate a timeline of keystrokes, examine spans of related edits (e.g., drafting a sentence or clarifying an argument), and assign one or more intention labels from a predefined taxonomy via a dropdown menu. This design allows annotators to interpret not only what changes were made, but why they occurred, linking textual behavior to underlying cognitive intent.
@@ 3.3 Writing Intention Annotation
| Planning | Idea Generation | Formulate and record initial thoughts and concepts. | writing down keywords of a paragraph beforehand (e.g., “..%[Comment out] main point: artifacts lack in human subjectivity..”) | 7.0% |
@@
| Revision | Fluency | Fix grammatical or syntactic errors in the text or LaTeX commands. | “We desiginingned several experiment setups for the LLM evaluations as described..” | 1.4% |
| Coherence | Logically link (1) multiple sentences within the same paragraph; (2) any two subsequent paragraphs; or (3) objects to be consistent. | “Each comment was annotated by three different annotators , which and we achieved high inter-annotator agreement.” | 3.3% | |
@@
| Structural | Improve the flow of information by modifying the location of texts and objects. | “..human alignment to compare the alignment between lexicon-based preferences and humans’ original preferences. First, we calculated the score from each human participant to compare the alignment between..” | 3.7% | |
@@
Building on cognitive theories of writing (Flower and Hayes, 1981) and recent empirical studies on scholarly revision (Du et al., 2022b; Koo et al., 2023), we developed a taxonomy of 15 distinct writing intentions grouped into three overarching categories (Table 1). Following Pustejovsky et al. (2017), two annotators conducted iterative open coding on initial subsets of keystrokes to inductively identify recurring cognitive patterns. Each span of edits was treated as a meaningful writing unit, such as composing a sentence or improving phrasing, and assigned an appropriate intention label. Disagreements were resolved through multiple rounds of discussion with a cognitive linguist, who refined label definitions and boundaries to ensure coherence.
To validate the taxonomy, we applied two principles from taxonomy design (Nickerson et al., 2013; Kundisch et al., 2021): mutual exclusivity (each edit corresponds to one primary intention) and collective comprehensiveness (taxonomy covers all observed behaviors). After iterative refinement, we achieved a weighted F1 inter-annotator agreement of on a 1K-keystroke subset, confirming strong reliability. See details in Appendix B.
Post-processing
After annotation, we post-processed the data to improve usability and ensure privacy. All personal identifiers were removed, and keystrokes labeled as non-informative artifacts (e.g., scrolling, idle typing) were filtered out. Remaining entries were segmented into discrete intention spans suitable for analysis and model training. (See details in Appendix A.3).
4 ScholaWrite
ScholaWrite represents the first large-scale publicly available dataset capturing the cognitive dynamics of end-to-end scholarly writing. It links every keystroke to the author’s underlying writing intention, allowing direct observation of how ideas evolve into finalized text. Each annotated entry follows a standardized schema connecting text changes, contextual metadata, and intention labels, as shown below, illustrating a Fluency edit, in which the writer corrects “expct” to “aspect.”
4.1 Overview
The final ScholaWrite dataset contains 61,504 keystroke-based writing actions collected across five Overleaf projects, each of which culminated in an arXiv preprint authored by graduate researchers. Every action is annotated with a cognitive writing intention drawn from the 15-category taxonomy described in Section 3.3. Table 2 summarizes key statistics across projects. In total, the dataset captures over 62K text edits, encompassing more than 118K words added or deleted across multi-month writing trajectories. Each project contains thousands of temporally ordered keystrokes, providing a detailed chronicle of the writing process at the sentence, paragraph, and document levels.
4.2 Analytical Units: Capturing Writing Flow
Because the dataset records precise timestamps in milliseconds, we can reconstruct the temporal structure of scholarly writing. To analyze this structure, we define two complementary units: intention sessions and sessions.
Intention session.
An intention session is a continuous period during which the writer focuses on a single cognitive goal, such as producing text, improving clarity, or inserting a figure, before switching to another intention. These sessions typically last from a few seconds to several minutes, capturing a concentrated burst of focused cognitive activity. We apply two constraints for meaningful segmentation: (1) inactivity longer than 10 minutes marks the end of a session; (2) sessions shorter than 30 seconds are discarded to exclude trivial edits. This definition allows us to quantify how long writers sustain specific cognitive modes before transitioning to new tasks.
| Project | 1 | 2 | 3 | 4 | 5 |
| # Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
| # Keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
| # Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
| # Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
Session.
In contrast, a session aggregates all writing actions within a continuous working period, regardless of how many intentions are involved. It represents the natural rhythm of scholarly writing, how authors weave together planning, drafting, and revising while working on a particular section or idea. Sessions end after 10 minutes of inactivity and must last at least one minute.
5 Analyses
ScholaWrite offers an unprecedented window into the dynamic, cognitive process of scholarly writing. Our analysis pursues two complementary goals: (1) to uncover how scholarly writing intentions evolve, interact, and shift across time, and (2) to inform the design of cognitively aligned writing assistants capable of supporting real-world research workflows. We focus on three research questions:
5.1 Which Writing Intentions Dominate?
We first examine which of the fifteen annotated writing intentions account for the greatest share of time and cognitive effort.
| Intention | % Time | Intention | % Time | ||
| Idea Generation | 4.8 | Fluency | 3.3 | ||
| Idea Organization | 0.6 | Coherence | 6.9 | ||
| Section Planning | 2.6 | Clarity | 13.4 | ||
| Text Production | 40.3 | Structural | 4.0 | ||
| Object Insertion | 8.4 | Linguistic Style | 2.6 | ||
| Citation Integration | 2.4 | Scientific Accuracy | 1.2 | ||
| Cross-reference | 2.4 | Visual Formatting | 6.1 | ||
| Macro Insertion | 0.8 |
Time Distribution across Intentions
Table 3 summarizes the average share of time spent per intention across five projects. Text Production dominates with 40.3% of total time, followed by Revision intentions – mainly Clarity, Coherence, and Visual Formatting – at 26.4%. In contrast, Planning activities (e.g., Idea Generation, Organization) take less than 5%.
These results suggest that scholars devote most of their time to articulating and refining text rather than explicit planning. The considerable effort spent on revision and formatting underscores that much of scholarly writing involves clarifying and visually shaping ideas, an aspect often overlooked in computational models of writing.
Session Duration and Cognitive Effort
We analyze the average duration of each intention session – a continuous period focused on one cognitive goal. Table 4 shows that Idea Organization sessions last the longest (3.7 min), followed by Object Insertion (2.9 min) and Visual Formatting (2.2 min), all showing high variability due to complex tasks like reorganizing structure or managing visuals.
In contrast, Text Production, though dominant in total time, consists of short, frequent bursts (mean 2.2 min). This contrast reveals two cognitive modes: (1) short, frequent drafting episodes, and (2) long, focused planning or visual composition that require sustained attention.
| Intention | Mean | SD | Max | Count |
| Idea Generation | 1.87 | 1.94 | 11.40 | 76 |
| Idea Organization | 3.71 | 4.58 | 13.79 | 7 |
| Section Planning | 1.37 | 1.57 | 9.36 | 38 |
| Text Production | 2.20 | 2.53 | 28.14 | 724 |
| Object Insertion | 2.88 | 3.05 | 21.61 | 108 |
| Citation Integration | 1.70 | 1.53 | 7.53 | 34 |
| Cross-reference | 1.08 | 0.77 | 3.98 | 38 |
| Macro Insertion | 1.43 | 1.05 | 3.71 | 7 |
| Fluency | 0.66 | 0.18 | 1.06 | 10 |
| Coherence | 1.46 | 1.71 | 9.62 | 55 |
| Clarity | 1.64 | 1.61 | 9.53 | 243 |
| Structural | 1.50 | 1.43 | 8.42 | 53 |
| Linguistic Style | 2.05 | 2.55 | 10.81 | 24 |
| Scientific Accuracy | 1.97 | 1.60 | 6.44 | 20 |
| Visual Formatting | 2.22 | 2.66 | 14.25 | 87 |
5.2 How Are Writing Intentions Intertwined?
Scholarly writing rarely unfolds linearly. Instead, it involves recursive alternations between complementary intentions, writers generate ideas, structure them, produce text, and immediately refine it. As seen in Figure 1 (and other projects in Appendix Figure 14), the intertwined processes of writing reveals dense bidirectional flows across intentions.
Transition Patterns
The transition probability matrix (Figure 2) shows that more than half of all Planning actions (e.g., Idea Generation) are immediately followed by Text Production. Conversely, Idea Organization often leads back to Idea Generation, suggesting that structuring ideas can prompt new content creation. Within Implementation, most transitions also return to Text Production (>30%), except for Macro Insertion, which frequently precedes Idea Generation. In Revision, strong recursive loops emerge, especially among Clarity, Fluency, and Coherence, highlighting how writers iteratively refine language and argumentation.
Overall, the results indicate that planning, implementation, and revision are not distinct stages but interdependent cycles continually revisited throughout the writing process.
Multitasking and Alternation Patterns
Writing sessions are highly multitasking in nature. As shown in Figure 3, 57% of all sessions involve three or more distinct intentions, reflecting frequent switching between cognitive modes within a single sitting. This dynamic interplay reveals how writers manage multiple goals in parallel.
Frequent 6-gram intention sequences (Appendix Table 9) confirm this recursive pattern. The most common alternating pairs include: (Text Production, Clarity), (Text Production, Object Insertion), and (Idea Generation, Idea Organization). These short feedback loops illustrate the micro-structure of scientific writing: writers repeatedly produce, adjust, and reorganize – blending idea formation with textual execution.
| Intention | Early | Late | Trend |
| Idea Gen. | 1.6 (23) | 2.5 (11) | longer, fewer |
| Idea Org. | 2.6 (2) | — | early-only |
| Sect. Plan. | 1.6 (13) | 1.0 (17) | more, shorter |
| Text Prod. | 2.1 (221) | 2.0 (227) | stable |
| Obj. Insert. | 2.4 (28) | 3.4 (42) | up both |
| Cite Int. | 1.4 (16) | 1.8 (6) | fewer later |
| Cross-ref. | 0.7 (8) | 1.2 (18) | more later |
| Clarity | 1.5 (43) | 1.3 (99) | more, quicker |
| Structural | 2.4 (8) | 1.4 (29) | more later |
| Ling. Style | 3.5 (3) | 1.7 (12) | concise edits |
| Vis. Form. | 3.6 (16) | 1.7 (43) | quick fixes |
5.3 Are writing patterns stable throughout the process, or do they change across phases?
Scholarly writing is not static, it evolves as authors move from idea formation to refinement. Beyond the aggregate trends observed earlier, we examine how writing intentions shift across time by comparing the early phase (first third of project duration) with the late phase (final third).
Which intents dominate at different stages?
Figure 4(a) shows how writing time is allocated across phases. Planning intentions are more frequent early (9.1% vs. 5.4%), reflecting front-loaded structuring. Within Implementation, Citation Integration is higher early, while Cross-referencing increases late as authors add visuals and refine internal links. Revision grows sharply in the late phase (46.7% vs. 29.6%), consistent with end-stage polishing for coherence and presentation.
Table 5 summarizes session-level differences. In Planning, Idea Generation and Organization sessions decrease in count but lengthen, suggesting deeper cognitive engagement later. Section Planning rises as writers reorganize material for submission. For Revision session counts increase substantially, though durations shorten, indicating rapid, localized edits characteristic of final revisions.
How are intentions intertwined across stages?
Figure 15 in Appendix compares transition probabilities across phases. Early writing features strong flows into Implementation, whereas late writing shifts toward Revision, marking the transition from building content to refining it. Some intention paths also evolve — for example, transitions from Macro Insertion lead to Idea Generation early but to Section Planning late, as layout decisions become more prominent. Writers also juggle more intentions later in the process. As shown in Figure 4(b), 19.4% of late sessions include over five distinct intentions (vs. 14.8% early), reflecting increased multitasking and integration of planning, drafting, and revising.